0 0
Add a comment Improve this question Transcribed image text
Answer #1

-\frac{\bar{h}^2}{2m}\frac{d^2\Psi }{dx^2}=E\Psi

\Rightarrow \frac{d^2\Psi }{dx^2}=-\frac{2mE}{\bar{h}^2}\Psi

This equation is of the form,

\frac{d^2y }{dx^2}=-k^2y

And its soluton is given by,

y=Asin(kx)+Bcos(kx)

In this case,

k^2=\frac{2mE}{\bar{h}^2}

\Rightarrow k=\frac{\sqrt{2mE}}{\bar{h}}

So, we have our solution,

\Psi(x) =Asin(kx)+Bcos(kx)

\Psi (0)=0

so we have,

\Psi(0) =Asin(k*0)+Bcos(k*0)

\Rightarrow B=0

\Psi(x) =Asin(kx)+Bcos(kx)

\Rightarrow {\Psi}'(x) =Akcos(kx)-Bksin(kx)

{\Psi }'(0)=1

So, from this we have,

\Rightarrow {\Psi}'(0) =Akcos(k*0)-Bksin(k*0)

\Rightarrow 1=Ak

\Rightarrow A=\frac{1}{k}

So, our wave equation is given by,

\Psi(x) =\frac{1}{k}sin(kx)\; where\;k=\frac{\sqrt{2mE}}{\bar{h}}

Or putting the value of k,

\Psi(x) =\frac{\bar{h}}{\sqrt{2mE}}sin(\frac{\sqrt{2mE}}{\bar{h}}x)

b)

For spatial period, that is the wavelength, we find the distance between the two alternate points where this function vanishes

Wave Wavelength Wavelength Distance

So,

\Psi (x)=0

\Rightarrow \frac{\bar{h}}{\sqrt{2mE}}sin(\frac{\sqrt{2mE}}{\bar{h}}x)=0

\Rightarrow sin(\frac{\sqrt{2mE}}{\bar{h}}x)=0

\Rightarrow \frac{\sqrt{2mE}}{\bar{h}}x=\pm n\pi

Distance between alternate point = 2*pi,

So,

\frac{\sqrt{2mE}}{\bar{h}}\lambda =2\pi

\Rightarrow \lambda =\frac{2\pi\bar{h}}{\sqrt{2mE}}

c)

Given that,

E=\frac{p^2}{2m}

So, we get,

\Rightarrow \lambda =\frac{2\pi\bar{h}}{\sqrt{2m*\frac{p^2}{2m}}}

\Rightarrow \lambda =\frac{2\pi\bar{h}}{p}